Two trains going in opposite directions leave at the same time. Train B travels 25 mph faster than train A. In 4 hours the trains are 460 miles apart. Find the speed of each.
Answer by addingup(3677)

distance = speed*time
460 = (x+x+25)*4 Divide both sides by 4
460/4 = 2x+25
115 = 2x+25 Subtract 25, both sides
90 = 2x Divide both sides by 2 and flip so it looks nicer
x = 45 is train A. And train B:
45+25 = 70
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Quick check:
Train A: 45
Train B: 70
Total... 115 is the combined speed at which they are traveling away from each other.
460 miles/115mph = 4 hours. We have the correct answer.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Re. "looks nicer": x=45 is the same as 45=x but convention says we should have the unknown on the left.
What is convention?:a way in which something is usually done, especially within a particular area or activity
Good luck,
J